A Complete Bibliography of Publications in Communications in Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

(0, 1) [PR75b]. 2 × 2 [SO73]. 2 × 2 × 2 [GH75]. 2 × C [SO73, SO74b]. A [PF73]. α_{3} [Bur73a]. α_{4} [Bur73a]. B [PF73]. C [KS74b]. C − 1 [SO73]. χ^{2} [Kim74, SO73]. c × c [HM73]. D [Ban74, Ago73]. E(y) = α + δx + βρ^{x} [BW74]. k [Dud74, PH73, ST75, SO74b]. L_{vc} [Nag75]. n = 100 [Ago73]. p [BR74b, Nai75, TC74]. q [BR74b]. Σ = I [Mui74]. t [Ahm75a, Ano75m, Ima74, SO74a]. T^{2} [GG75]. U [FR75, PP74]. w′ [Ago73]. X_{1} + X_{2} [Loc75]. X_{1}/X_{2} [Loc75].

-class [ST75]. -dimensional [TC74]. -matrix [PR75b, KS74b]. -optimality [Ban74]. -ratio [PP74]. -sample [PH73]. -shaped [FR75]. -statistic [Ima74, SO74a]. -statistics [SO73]. -variates [GG75].

accelerated [Sin74]. acceptable [Joh74]. act [PF73]. action [SB75b]. Adaptive [Fri75, RH73]. addition [KLR73]. additional [DT73]. adjustment [MM74]. Admissible [Lin74a]. after [MH73]. against [PH73]. algorithm [CR75, MW74]. allocation [SS74]. alternative [Beh75, SN73]. alternatives [Lep75, PH73]. altoregressive [Pag73]. among [CL74]. analogues [Kas74]. analysing [OM73]. Analysis [Kas74, Adi75, AB75, BHOO74, CH74, FF75, GW75, MW74, PC74, SO75, Sug73, TLD75, dW74]. analytic [Saw75b]. ancillary [Fu74]. Anderson [PR75b]. application [Bec73, GD74, MY75, SC74c, TS74, TLD75]. applications [GG74, Saw73, Ste74, Sug73]. approach [Hoc75a, SO75, War74]. Approximate [SO73, Bur73b]. approximation [GH75, Ima74, MP74, Phi73]. Approximations [PC75, Bur73c]. arbitrary [PH73, Pat75, RY74]. arising [Sin74]. assay [PC75]. associated [KA74b, UWH73]. Asymptotic [GCK75, Sch73a, SN73, WS74, HC74]. Asymptotically [Hen75, Lep75, Ahm75a]. attracted [Roh73]. autocorrelated [HL73]. autocorrelation [Pie75]. autocorrelations [Rob74]. automatic [WW75a]. autoregressive [MT74, Vin73]. average [Cla74].

bad [Rya74]. balanced [PF73, Sah74]. bands [Per74]. based [GM73, GH74, Wei73]. basic [Sen74]. Bayes [BS75, Haa75, PC75, Sah74, SB74, SB75b, dW74]. Bayesian [BB75, Bro74, PC74, War74]. been [MH73]. Behrens [Hoc75a, SS75]. Bernoulli [BR74a, Bam75, Cro75]. best [Ano75m]. between [Bur73b, GM73]. bias [KS75a]. binomial [BK73, BB75, Bur73b, Bur73c, Con75, Ham74a, Joh74]. birameters [Wol73]. bivariate [Ahs74, Ham74a, KS75b, KS75c, TS74]. block [Pat75]. blocking [TR74]. Board [Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l]. bound [Lwi74, Saw74]. Bounds [KR74]. Buffon [GD74]. busy [Jai75].

calculation [Ham74b]. calibration [TM73]. Carlo [Ago73, HL73]. case [GWS73, ND75, Phi73]. cases [Eic74]. categorical [Kas74]. categories [Dev75, HC74]. cell [GH75]. censored [BK73, KD75, LHS74]. censorship [PH73, Pat75]. central [Rub74b, Rub74a]. centroid [Kim74]. Certain [Jan75, And75a, BOL74, LSR73, RY75]. Certainty [Lin74b]. chance [HM73, Lin74b]. change [BJ75]. changing [Bro74]. characteristic [Sug73]. Characterization [Gup74, Ahs74, AK74, Con75, Gup75a, KJ74, Rao74]. Characterizations [Gal75, Gup75b]. Characterizing [FR75, Kor75]. chi [Cam75, Ham74a, Ham74b, Kat73, RR74, Rub74b, Rub74a]. chi-square [Cam75, RR74, Rub74b, Rub74a]. chi-squared [Ham74a, Ham74b]. class [Ahm75a, ST75, WS74]. classifiers [Dev75]. clinical [DK74, HW75]. Close [Bur73c]. closeness [Zac74]. cluster [CH74]. clustering [Sch73b]. coefficients [Woo74]. column [SK73b]. combination [KW73]. combinations [BR74b]. Comment [Haw77, Tik77]. Comparing [WC74, BB75, Nai75, SB74]. comparing-tail [SB74]. comparison [Ago73, Cla74, CR75, OC74, Sin73, SA74, SS75, WS74]. comparisons [BG75, Dav73, PS74b]. Complete [PP73]. completely [WW75a]. complex [ND75]. component [Lot73]. components [Ala74, BS75, Che74b, Joh73, Sah74, SB75b]. Composite [Gra73]. compromise [Sha75b]. computer [LM73]. concepts [Kas74]. concordance [Woo74]. conditional [And75a, Gup75a]. conditions [Hoc75a]. Confidence [Cro75, Rea74, AA74, BR74b, Bur73c, Dun74, Per74, SS75]. confounded [KS74b]. conjugate [Fri75]. Connectedness [SK73b]. consistency [KV75]. constant [GP75]. constraint [Lin74b]. constraints [War74]. constructing [Hig75, MH73, OC74]. construction [RN74]. containing [Rya74]. contiguous [Lep75]. contingency [GH75, Ham74a, HC74, MY75, SO73, SO74b]. continuous [CS74]. contrasts [VR73]. contribution [Jai75]. convexity [Coo74]. Cornish [MG75]. correctness [PS74a]. correlated [Dev75, KW73, WC74]. correlation [BOL74, Ham74a]. covariance [Adi75, PP74, Rob74, RY74, RY75, RL75]. covariance-time [RL75]. credibility [Rea74]. criteria [KS75a, SA74]. criterion [ND75, Nag75, OC74]. cross [WW75a, WW75b]. cumulants [MG75]. curve [BW74, Cha75, SK73a, WW75a]. curves [RL75, TLD75]. cycle [Gra73].

Darmois [KLR73]. Data [DK74, Kas74, KD75, LHS74, OM73, SO75, WC74]. defined [KK74]. degree [NB73, PR75b, WW75b]. degrees [Rub74a]. dendrite [CH74]. densities [Hig75]. Density [Car75, KV75, SO74a, Van73, WW75b]. dependence [Loc75]. dependent [BR74a]. Derivatives [Sug73]. derived [BR74a, SO73]. Design [GA75, KS75a, HW75, MH73, MM74]. designs [Bec73, Gra73, Haz73, KS74b, MW74, OC74, PF73, Pat75, RN74, Rak74, SK73b, SD75, TM73]. detection [BJ75]. determinant [SA74]. Determinants [GCK75, Ste74]. determining [WW75b]. deviation [Roh73]. different [CR75]. digital [MP74]. dimensional [TC74]. direction [Fri75]. discrete [Gup74, GG74, Gup75b, KK74, Kor75]. discriminant [LSR73]. dispersion [Cha75, DS75]. distributed [Lin74b, Phi73]. Distribution [GG75, HS75, Ahm75a, Ahs74, AK74, AR75, And75a, BR74a, FR75, GH75, Gup75a, Ima74, KA74a, Kat73, KD75, KW73, sL75, MSF73, ND75, PC75, PR74, PS74b, RH73, Rao74, SC74a, SC74c, Saw73, Saw74, Sin75, SG75, SB75a, SO73, Sun74, SB75b, TC74, Wan73, WK73, Woo74, WG75]. distribution-free [Ahm75a, RH73]. Distributions [GG74, GCK75, Bli74, BR74b, BS75, Bur73a, Bur73b, CF75, Con75, GP75, Gup74, Gup75b, Joh74, JH74, JK75, KK74, KS75b, KS75c, Kor75, KJ74, Que75, SO74a, Sen74, Sha75a, SN73, Zac74]. divergence [CL74]. Does [DS75]. dosage [Eic74]. double [ST75]. drift [BJ75]. Durbin [Vin73].

Editorial [Owe73a, Owe73b, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l]. Effects [HL73, BOL74, BG75, MP74, PR75b, Sah74]. efficiency [HC74, Lit74]. Efficient [MT74, Hen75]. elimination [RN74]. ellipsoids [GP75]. Empirical [LM73, BS75, Haa75, SB75b]. entry [GH75]. environmental [Rob74]. equal [sL75]. equal-mean [sL75]. equality [Cho75, PP74]. equicorrelated [HS75]. equivalent [Lin74b]. error [Lin74a, OC74, TC74, YJB74]. errors [HL73]. estimate [HS74, Lwi74, PC75, Rob74, dW74]. estimates [Dav74, Joh74, Sug74]. estimating [Hoc75b, MT74, PR75a, Zac74]. Estimation [Ala74, Dud74, KA74a, KS74a, KD75, RL75, Rya74, Sha75b, Sin74, UWH73, Bli74, Blu75, BS75, Car75, Cha75, Cla74, DD74, Gra73, GA75, Hen75, KS75a, KV75, MY75, RY74, Sen74, Sie73, Sli75, Sun74, Van73, WW75b, Wan73, WG75, Bli74]. estimator [KV75]. Estimators [Bam75, BK73, AR75, AA73, EB73, GWS73, HL73, Lin74a, PP73, Sah74, ST75, Tan74, WK73]. Evaluation [CS74, Ste74]. even [Rub74a]. Exact [BHOO74, DT73, Kat73, ND75, Ima74]. example [Ban74]. exchangeable [Sha75a]. expansions [DT73, MG75]. expectation [Gup75a]. expected [KR74, Zac74]. experimental [MW74]. experiments [KS74b, OM73]. exploiting [Rob74]. Exponential [Phi73, DD74, KJ74, PC74, PC75, RR74, SB75a, SB75b, TS74, Zac74]. exponentiality [TRM74, Tik74b]. exponentials [BK73]. Extension [KLR73, Cam75, GD74, PR75b, SK73a, SO74b, TS74]. extreme [MSF73]. extreme-value [MSF73].

factorial [KK74, KS74b, OM73, Rak74]. failure [Ahm75b, KD75]. failures [HW75]. families [KS75b]. family [DD74, KK74, RR74]. Farlie [JK75, Sha75a]. Federer [PR75b]. Filter [MM74]. financial [PS74a]. Finite [Tho73, CS74]. first [MT74]. first-order [MT74]. Fisher [Hoc75a, MG75, SS75, Tan73, TS74]. fit [LHS74, MSF73, RR74]. fitting [SO75, SK73a, WW75a]. Fixed [AA74, Per74, HW75, KS75b, SN73]. folded [Sun74]. form [Saw75b]. formal [Sah74]. forms [KW73]. formulas [DT73, SN73]. formulation [Rak74]. foundation [Tho73]. fractional [Rak74]. framework [CS74, MW74]. free [Ahm75a, RH73, Wol73]. freedom [Rub74a]. French [WW75a]. Friedman [Pat75]. function [Ahm75b, KA74a, LSR73, Phi73, Wol73]. function-free [Wol73]. functions [Bam75, FR75, Ham74b, KA74a, KLR73, WW75a].

gamma [AR75, Bli74, Rub74b, WG75]. Gaussian [CF75]. general [Bur73a]. Generalization [Vin73]. Generalizations [SC74b]. generalized [Gho74, Jai75, JK75, Kle73, MW74, PH73, Pat75, Sha75a, ST75]. generate [JH74]. generation [LM73]. genetic [Rob74]. geometric [BHOO74]. Geometrical [Hig75, Rak74]. GMANOVA [Kab75]. goodness [LHS74, MSF73, RR74]. goodness-of-fit [MSF73, RR74]. gradually [Wei73]. grid [Bur73a]. grouped [BK73]. grouped-censored [BK73]. growth [Cha75, TLD75]. Gumbel [JK75, Sha75a].

Haenszel [SO74b]. Hermitian [Sug73]. heterogeneity [RN74]. heteroscedastic [Sin74]. higher [PR75b, Vin73]. histogram [KV75, Van73]. Hogg [Beh75]. Hotelling [Ban74, GG75, PS74b]. Hyperbolic [GM73]. hypergeometric [Bur73b, Jan75]. hypotheses [Phi73, SC74a, Wol73]. hypothesis [PY73].

identically [Phi73]. if [Lin74b]. II [Bli74, SC74a, SC74b]. important [DT73]. Improved [AR75, Saw75b]. incorporate [BG75]. increasing [Wei73]. Independence [Arn73, SN73]. independent [GG75, Phi73]. inequality [War74]. Inference [BR74a, And75b, Eas75, Hoc75a, Jan75, Tan73, Tho73]. inferences [Bro74]. inferential [Hig75]. information [AB75, GA75, HS74, Sli75]. initial [MH73]. input [MH73]. integral [Cam75, Rub74a]. integrated [OC74]. intensity [Saw75a]. intervals [AA74, BR74b, Dun74]. intraclass [BOL74, RY74]. inverse [CF75, MW74, SR73, TM73]. investigation [LM73]. involving [BS75, Cho75]. iterative [Sug74].

jackknife [And75b, GWS73]. Jacobian [Saw75b]. Jacobians [Saw73]. join [Sin74]. Joint [BR74b]. Jonckheere [PH73].

kind [Sin75]. kinds [SC74c]. knowledge [AA73]. known [Kas74]. Kolmogorov [Kim74, Lit74]. Kruskal [HC74].

Lagrangian [Con75]. Large [And75b, DK74, Roh73]. largest [Blu75, HS75, Mui74]. latent [Mui74, Saw74]. law [Roh73]. least [And75b, Che74a, HL73, Lot73, Rao73]. least-squares [Lot73]. level [MT74]. levels [Bur73c]. life [BK73, Sin74, Sli75, SB75a]. life-testing [SB75a]. likelihood [AR75, Cha75, Dav74, Dav73, EB73, Phi73, PP74, RY75, Sie73, Sug74, Tan74]. limits [Bur73c, Cro75, TC74]. Linear [Dev75, AB75, BR74b, Eic74, GA75, GM73, KA74a, KS74a, Kle73, Kor75, KW73, LSR73, MP74, MB74, Per74, RY75, Sin73, UWH73, War74]. linear.hypotheses [Kle73]. linearity [Adi74]. lines [Dav74]. location [Bli74, EB73, KA74a, KS74a, Lep75, Sha75b]. location-scale [EB73]. logarithmic [PR74]. logit [SO74b, Sug74]. lognormality [KAR75]. loss [Lin74a]. lower [Lwi74, Saw74].

main [PR75b]. MANOVA [FF75, TLD75]. Mantel [SO74b]. marginals [KS75b]. Markov [DS75, KA74b]. Markovian [Sch73a]. match [Bur73a]. Matrices [GCK75, Cha75, JH74, PP74]. matrix [DS75, KS74b, PR75b, RY74, RY75, SR73, SG75, ST75, Sug73]. max [PY73, PP74]. maximin [GH74]. Maximum [Cha75, Dav74, Sug74, AR75, Dav73, EB73, HW75, Sie73, SA74, SO73, Tan74]. Maximum-likelihood [Cha75]. mean [And75a, AA73, Blu75, sL75, Lin74a, MSF73, OC74, YJB74]. means [BR74b, Dud74, WC74]. measure [CL74, HS74]. measurement [CR75]. Median [Joh74, HC74]. Median-unbiased [Joh74]. method [CH74, Hig75, MH73, MT74, MY75, RN74, SK73a, Sug74, TM73, Van73]. methods [BHOO74, Fri75, Gel74, OC74]. minimax [Lin74a]. Minimum [WG75, KS75a]. missing [WC74]. mixed [KD75]. mixture [BK73, Dav74]. mixtures [Joh73]. modal [KV75]. model [BG75, Cha75, CR75, Dav73, GA75, GM73, Kab75, KLN75, PS74a, PC74, Sah74, Sin74, SO74b, Sug74, TS74, TLD75]. models [Jan75, KS74a, Kle73, MP74, Per74, Sen74, UWH73]. modified [Fri75]. Moment [DD74, Gup75b, ST75]. Moment-type [DD74]. moments [HS74, KK74]. monotonicity [Ahm75b]. Monte [Ago73, HL73]. Morgenstern [JK75, Sha75a]. most [BR74b, Joh74]. moving [Cla74]. multi [Tik74b]. multi-sample [Tik74b]. multicollinearity [GW75]. multinomial [Hoc75b]. multiparameter [WS74]. multiple [AA73, Dun74]. multivariate [And75a, BOL74, Bec73, CR75, DS75, FF75, Gel74, GP75, Hig75, Jan75, Kle73, Lin74a, PR74, Rao74, SS75, Sug73, WK73, WS74, dW74].

near [Dev75]. near-neighbor [Dev75]. needle [GD74]. negative [Joh74]. negative-binomial [Joh74]. neighbor [Dev75]. nested [Sah74]. Neyman [Jai75]. Nile [Tan73]. Non [Rub74b, Fu74, GA75, LSR73, Rub74a, SO74a]. Non-central [Rub74b, Rub74a]. non-linear [GA75]. non-normal [SO74a]. non-normality [LSR73]. non-uniqueness [Fu74]. noncentral [Cam75, JH74]. noncentrality [Ham74b, PR75a, dW74]. nonidentical [BS75, SB75b]. nonparametric [Ahm75b, Hen75]. nonregular [Bli74]. Normal [GH75, Ahs74, And75a, BOL74, Blu75, BR74b, BS75, CR75, DS75, Dud74, GWS73, HS75, Kim74, Lin74a, Rao74, RY74, RY75, Roh73, SC74a, SO74a, SB74, Sun74, WK73]. normality [Ago73, KS74a, KAR75, LSR73, Tik74a, Tik74b]. note [Che74a, sL75, MP74, NB73, PR74, RN74, Sha75a, Sin75, Tan74, TLD75]. nuisance [EB73, Tan74]. number [HW75, Wei73]. numerical [CR75].

observations [Adi74, Gal75, Gel74, PH73, Pat75]. Obtaining [MB74]. One [KS75b, VR73, Kat73]. One-and [VR73]. One-parameter [KS75b]. Optical [TM73]. optimal [BB74, Che74b, Eic74, Lot73, MH73]. optimality [Ban74, SA74]. optimum [KA74b, Lep75]. order [Arn73, BG75, Coo74, Gup74, KR74, KA74a, KLN75, LM73, MT74, RY74, ST75, Vin73, Wei73]. ordered [Dud74, HC74, PH73]. ordering [Car75]. original [Ban74]. orthant [Cho75]. orthogonal [And75b, Ham74b, VR73]. orthonormal [Ham74b]. Outlier [Tik77]. outliers [Haw77, Tik75]. overall [PS74a].

Pair [SD75, BG75]. Paired [Sin73, BG75, Dav73]. pairs [MB74]. parameter [AA74, Bli74, Ham74b, KS75b, MSF73, PR75a, Que75, Sli75, Wan73, dW74]. Parameters [Bur73a, BK73, Dav74, Dud74, EB73, GA75, Hen75, KA74a, KS74a, KD75, MSF73, Sha75b, Tan74, War74]. Pareto [AK74, Wan73]. partial [Saw74]. Partially [PF73]. pdf [GP75]. Pearson [Ham74a, Kim74]. Percentage [Nag75, ND75]. period [Jai75]. Periodic [WW75b]. petition [Shu74]. phase [GM73, SS74]. piecewise [GM73]. plantation [Rob74]. play [Rag74]. point [RL75, Sin74]. points [Car75, ND75, Nag75]. Poisson [Bur73b, Con75, Joh74, Saw75a]. polynomial [NB73]. polynomials [Ham74b]. population [BOL74, DS75]. populations [Ano75m, CS74, CL74, GH74, SB74, Tho73]. possibility [Rob74]. posterior [Lwi74]. power [Ago73, KA74a, PS74b, SB75a]. power-function [KA74a]. Powers [Mui74, Kim74]. prediction [Fri75, KLN75]. preference [MB74]. presence [EB73]. principal [Che74b, Lot73]. prior [AB75, AA73, GA75]. priori [Sli75]. probabilities [BB75, Cho75, Cro75, KA74b, Roh73, SB74, Zac74]. probability [Bam75, Cam75, Rub74a]. problem [Ban74, BS75, GD74, Haa75, Rak74, SB74, SS75, SB75b, Tan73]. problems [BK73, GW75, Jan75, MY75]. procedure [BB74, PR75b, SO74b]. Procedures [JH74, And75b, Ano75m, Cla74, GA75, GH74, Hig75, Wei73, Zac74]. process [Ala74, BJ75, KA74b, MT74, RL75, Saw75a]. processes [Nai75, Vin73]. product [Saw74]. programming [Sin73]. properties [BB74, Gal75, GP75, Gup75b, PF73, Sch73a]. property [Che74b, EB73, Gup74, KJ74]. proportional [Zac74]. proportions [Sch73a]. proposal [Beh75].

quadratic [KW73, LSR73]. quantal [PC75]. quasi [Con75]. quasi-binomial [Con75]. queues [Jai75]. quotient [GG75].

radial [TC74]. Randles [Beh75]. Random [GCK75, Bro74, FR75, JH74, KLR73, Lin74b, Roh73, Sah74, SG75]. Randomization [Eas75]. randomized [Hoc75b, Pat75, Sen74]. randomness [Sch73b]. range [AA74, Dud74]. Rank [Adi74, Lep75]. Ranking [Adi75, Dav73, Sin73]. ranks [Ima74]. Rao [Che74a]. Rapid [Gel74]. rate [Ahm75b]. ratio [PY73, PP74, RY75]. reducible [RY75]. reduction [DK74]. regimes [GM73]. region [SS75]. Regression [GW75, Adi74, AB75, BW74, Dav74, Eic74, GM73, Hen75, HKB75, Kor75, Lot73, MH73, MW74, Per74, PS74a, SK73a, Sin74, TM73, War74, YJB74]. Rejoinder [Tik77]. related [CF75, Hig75, MY75, WK73]. relation [GWS73, SO74b]. relations [Bur73b]. relative [Lit74]. Remarks [PR75b, PR75a]. renewal [KA74b, Sch73a]. Repeated [SC74a, SC74b, Adi74, SC74c]. repetitions [Rya74]. representations [KS75c]. response [Hoc75b, OC74, Sen74]. restricted [AA74]. restrictions [Lot73, YJB74]. result [Cam75, Rub74a]. results [Kab75]. revisitation [UWH73]. revisited [Rub74b]. Ridge [HKB75]. right [PH73, Pat75]. Robustness [LSR73]. root [Mui74, SA74, Sug73]. roots [Saw74]. rotation [Gra73]. row [SK73b]. row-column [SK73b]. rows [Rob74]. Ruben [Cam75]. rule [Rag74]. rules [Nai75].

Sample [SS74, And75b, Gra73, Hoc75a, PH73, Sch73a, Shu74, Tik74b, Zac74]. samples [Que75, Rya74]. sampling [BOL74, CS74, GH74, Rag74, SS74, Sie73]. satisfying [KLR73]. scale [EB73, Hen75, KA74a, Lep75, Sha75b, Wan73]. scales [CR75]. scaling [Sug74]. scheme [Pag73]. schemes [Hoc75b]. score [Adi74]. scores [KA74b, MB74]. search [Eic74]. seasonal [MM74]. second [Coo74, NB73, Sin75, ST75]. second-order [Coo74]. selected [MH73]. Selection [Ano75m, BB74, FF75, GH74, Nai75]. sequence [Bam75, Bro74]. Sequential [BJ75, Blu75, Eic74, Sie73, Wan73, Zac74]. seriation [Gel74]. series [DT73, Fri75, MM74, PR74]. set [HS75, Lot73, MH73]. several [CL74, GH74, LM73, PP74]. Shannon [HS74]. shaped [FR75]. signed [Ima74]. significance [BOL74, SC74a, SC74b, SC74c]. significant [HM73]. similar [Ahm75a]. similarities [Gel74]. Simple [War74, Joh73, PC75]. simplex [MY75]. simulations [HKB75]. Simultaneous [Hoc75a, VR73]. singular [Haz73, Saw73, Saw75b]. situations [Tik74b]. size [SS74, Zac74]. Skitovic [KLR73]. small [And75b, Cro75]. smaller [Gal75]. Smirnov [Kim74, Lit74]. Smirnov-type [Lit74]. smooth [GM73]. smoothing [WW75b]. solution [SB74]. solutions [SS75]. Some [AA73, Bur73b, GP75, GCK75, Gup75b, HS74, Joh73, Kab75, Nai75, PR75a, RY75, Sah74, CS74, DT73, Hig75, HKB75, JK75, Kas74, Kor75, Lit74, Ste74, SS75]. space [GD74]. specifying [FR75]. spectral [WW75b]. sphericity [ND75]. spline [WW75a]. splines [WW75b]. square [Cam75, Kat73, OC74, RR74, Rub74b, Rub74a, SO75, YJB74]. squared [Arn73, Ham74a, Ham74b, Lin74a]. squares [And75b, Che74a, HL73, Lot73, Rao73]. stage [AA73, Hoc75b, Sah74]. standard [Nai75]. state [KA74b]. statements [PS74a]. states [KA74b]. Stationary [KA74b, Fri75, MT74, Pag73, RL75]. Statistic [Tik77, GWS73, Haw77, Ima74, RR74, SO74a, Tik74a, TRM74, Tik75, Vin73]. Statistical [CF75, Wei73, Eas75, Saw73]. statistics [Ahm75a, Arn73, Fu74, Gup74, KR74, KA74a, KLN75, LM73, NB73, PP73, SO73, SN73, Wei73]. Steck [Ste74]. Stirling [Sin75]. stochastic [YJB74]. strategies [CS74]. strategy [GH74, TR74]. Stringency [Haa75]. structure [Haz73, RY74, Sch73a]. structures [DS75, RY75]. Student [Ahm75a]. study [HL73]. subject [PH73, Pat75]. success [Bam75]. sufficient [PP73]. sums [FR75, GG74, Roh73]. surface [Coo74, OC74]. surveillance [PC74]. survey [Shu74]. Suspected [Tik77, Haw77, Tik75]. symmetric [Sug73]. symmetrical [Joh73, KS74b]. symmetry [Lit74]. system [Bur73a]. table [GH75, Kat73, SO73]. tables [Ham74a, HM73, HC74, MY75, SO75, SO74b]. tail [Car75, SB74, Zac74]. tailed [VR73]. techniques [LM73]. terms [Bur73b, DT73, Gal75]. test [Ahm75b, Beh75, Coo74, HC74, Ima74, Kat73, Kim74, LHS74, MSF73, Mui74, PH73, Pat75, PS74a, PY73, SC74a, YJB74]. Testing [Kle73, RY74, Tik74b, Tik77, BK73, Haw77, HM73, Phi73, Sli75, SB75a, SN73, Sun74, Tik74a, TRM74, Tik75, Wol73]. Tests [KAR75, Saw75a, Sch73b, Adi74, Adi75, Ago73, BOL74, Dun74, Joh73, Lep75, Lit74, PP74, RH73, RR74, RY75, SC74b, SC74c, Sin74, WS74]. their [FR75, Gup75b, KK74, SC74c]. theorem [KLR73]. theory [Che74a, Rao73, Saw73]. three [Sah74]. three-stage [Sah74]. time [BR74a, Fri75, KD75, KJ74, MM74, RL75]. Tolerance [TC74, PC75]. toxicity [Eic74]. trace [PY73, PS74b, SA74, SN73]. trace-ratio [PY73]. transformation [Saw75b]. transformations [Saw73]. transformed [Woo74]. Transforming [Que75]. transition [GM73]. translated [SB75b]. trend [Pie75]. trial [HW75]. trials [BR74a, Bam75, Cro75, DK74]. truncated [GG74]. truncation [Que75]. Two [Hoc75b, SC74c, BHOO74, Dav74, Gal75, Gra73, GM73, GG75, Hoc75a, KS74a, MSF73, RN74, SS74, SN73, Sug73, SB75b, VR73]. two-parameter [MSF73]. two-phase [GM73, SS74]. two-tailed [VR73]. two-way [BHOO74, RN74]. type [DD74, Lit74, Jai75]. types [LSR73].

UMVU [GWS73, WK73]. unbalanced [BHOO74]. unbiased [Joh74, WG75]. uncertain [Tan73]. underlying [And75a]. Unified [Rao73, Che74a]. Uniform [KV75, KS75c, SC74c, Sch73b]. uniformity [Que75]. uniformly [Lin74b]. unimodality [Gho74]. unique [Cha75]. uniqueness [Fu74]. unknown [Blu75, Car75, MSF73, MG75]. unsolved [Rak74]. untransformed [OM73]. urn [KLN75]. Use [Ima74, Sli75, AB75, GA75, Ham74b, WW75b]. used [YJB74]. useful [BK73]. using [AA73, CS74]. utilize [OC74].

validation [Shu74, WW75a, WW75b]. value [MSF73]. values [KR74, MH73]. Variable [FF75, CS74, Lin74b]. variables [Bro74, FR75, HS75, KLR73, Roh73]. variance [Ala74, And75a, Blu75, BHOO74, FF75, Lwi74, Rob74, RL75, Sah74, WG75]. variance-time [RL75]. variates [GG74, GG75, WC74]. various [HL73]. versus [KAR75]. via [KK74]. view [BHOO74]. viewpoint [BB75].

waiting [BR74a, KJ74]. waiting-time [BR74a]. Wallis [HC74]. Wanted [SR73]. Watson [Vin73]. way [BHOO74, Kat73, RN74]. Weibull [Bli74, MSF73]. weighing [Ban74, Bec73, Haz73]. Weighted [OM73, FR75]. well [Kas74]. well-known [Kas74]. which [OC74, Roh73]. width [AA74, Per74]. Wiener [Ala74, BJ75]. Wilcoxon [Ima74]. Wilk [Nag75]. winner [Rag74]. within [BG75, RR74, Rob74]. within-pair [BG75].

Yates [Ban74].

zero [Ham74a, Lit74].

References

Andrews:1973:SMS

[AA73] Robert L. Andrews and J. C. Arnold. Some multiple stage estimators of the mean using prior knowledge. Communications in Statistics, 1(6):553–560, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Andrews:1974:FWC

[AA74] R. L. Andrews and J. C. Arnold. Fixed width confidence intervals for restricted parameter range. Communications in Statistics, 3(2): 167–178, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anderson:1975:UPI

[AB75] R. L. Anderson and E. L. Battista. The use of prior information in linear regression analysis. Communications in Statistics, 4(6):497– 517, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Adichie:1974:RST

[Adi74] J. N. Adichie. Rank score tests for linearity of regression of repeated observations. Communications in Statistics, 3(7):671–678, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Adichie:1975:RAC

[Adi75] J. N. Adichie. Ranking in analysis of covariance tests. Communications in Statistics, 4(9):883–890, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Agostino:1973:MCP

[Ago73] Ralph B. D’ Agostino. Monte Carlo power comparison of the w′ and d tests of normality for n = 100. Communications in Statistics, 1(6):545–551, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ahmad:1975:CAD

[Ahm75a] Ibrahim A. Ahmad. A class of asymptotically distribution-free statistics similar to Student’s t. Communications in Statistics, 4 (9):863–871, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ahmad:1975:NTM

[Ahm75b] Ibrahim A. Ahmad. A nonparametric test for the monotonicity of a failure rate function. Communications in Statistics, 4(10):967– 974, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ahsanullah:1974:CBN

[Ahs74] M. Ahsanullah. A characterization of the bivariate normal distribution. Communications in Statistics, 3(3):259–262, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ahsanullah:1974:CPD

[AK74] M. Ahsanullah and A. B. M. Lutful Kabir. A characterization of the Pareto distribution. Communications in Statistics, 3(10):953– 957, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Alam:1974:EVC

[Ala74] Khursheed Alam. Estimation of variance components of a Wiener process. Communications in Statistics, 3(9):843–852, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anderson:1975:MVC

[And75a] Dallas W. Anderson. On the mean and variance of a certain conditional distribution when the underlying distribution is multivariate normal. Communications in Statistics, 4(12):1161–1166, 1975. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anderson:1975:LSJ

[And75b] Peter O. Anderson. Large sample and jackknife procedures for small sample orthogonal least squares inference. Communications in Statistics, 4(2):193–202, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1973:EBa

[Ano73a] Anonymous. Editorial Board. Communications in Statistics, 1(1): ebi, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1973:EBb

[Ano73b] Anonymous. Editorial Board. Communications in Statistics, 1(4): ebi, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1973:EBc

[Ano73c] Anonymous. Editorial Board. Communications in Statistics, 1(5): ebi, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1973:EBd

[Ano73d] Anonymous. Editorial Board. Communications in Statistics, 1(6): ebi, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1973:EBe

[Ano73e] Anonymous. Editorial Board. Communications in Statistics, 2(2): ebi, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1973:EBf

[Ano73f] Anonymous. Editorial Board. Communications in Statistics, 2(4): ebi, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1973:EBg

[Ano73g] Anonymous. Editorial Board. Communications in Statistics, 2(5): ebi, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1973:EBh

[Ano73h] Anonymous. Editorial Board. Communications in Statistics, 2(6): ebi, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBa

[Ano74a] Anonymous. Editorial Board. Communications in Statistics, 3(1): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBb

[Ano74b] Anonymous. Editorial Board. Communications in Statistics, 3(2): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBc

[Ano74c] Anonymous. Editorial Board. Communications in Statistics, 3(3): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBd

[Ano74d] Anonymous. Editorial Board. Communications in Statistics, 3(4): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBe

[Ano74e] Anonymous. Editorial Board. Communications in Statistics, 3(5): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBf

[Ano74f] Anonymous. Editorial Board. Communications in Statistics, 3(6): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBg

[Ano74g] Anonymous. Editorial Board. Communications in Statistics, 3(7): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBh

[Ano74h] Anonymous. Editorial Board. Communications in Statistics, 3(8): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBi

[Ano74i] Anonymous. Editorial Board. Communications in Statistics, 3(9): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBj

[Ano74j] Anonymous. Editorial Board. Communications in Statistics, 3(10): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBk

[Ano74k] Anonymous. Editorial Board. Communications in Statistics, 3(11): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1974:EBl

[Ano74l] Anonymous. Editorial Board. Communications in Statistics, 3(12): ebi, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBa

[Ano75a] Anonymous. Editorial Board. Communications in Statistics, 4(1): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBb

[Ano75b] Anonymous. Editorial Board. Communications in Statistics, 4(2): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBc

[Ano75c] Anonymous. Editorial Board. Communications in Statistics, 4(3): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBd

[Ano75d] Anonymous. Editorial Board. Communications in Statistics, 4(4): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBe

[Ano75e] Anonymous. Editorial Board. Communications in Statistics, 4(5): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBf

[Ano75f] Anonymous. Editorial Board. Communications in Statistics, 4(6): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBg

[Ano75g] Anonymous. Editorial Board. Communications in Statistics, 4(7): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBh

[Ano75h] Anonymous. Editorial Board. Communications in Statistics, 4(8): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBi

[Ano75i] Anonymous. Editorial Board. Communications in Statistics, 4(9): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBj

[Ano75j] Anonymous. Editorial Board. Communications in Statistics, 4(10): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBk

[Ano75k] Anonymous. Editorial Board. Communications in Statistics, 4(11): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:EBl

[Ano75l] Anonymous. Editorial Board. Communications in Statistics, 4(12): ebi, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anonymous:1975:SPB

[Ano75m] Anonymous. Selection procedures for the t best populations. Communications in Statistics, 4(11):987–1009, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Anderson:1975:IML

[AR75] C. W. Anderson and W. D. Ray. Improved maximum likelihood estimators for the gamma distribution. Communications in Statistics, 4(5):437–448, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Arnold:1973:ISO

[Arn73] Barry C. Arnold. Independence of squared order statistics. Communications in Statistics, 2(4):357–362, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Bamber:1975:EFP

[Bam75] Donald Bamber. Estimators for functions of the probability of success in a sequence of Bernoulli trials. Communications in Statistics, 4(10):941–953, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Banerjee:1974:OYO

[Ban74] K. S. Banerjee. On D-optimality of Yates’ original example in the Yates–Hotelling weighing problem. Communications in Statistics, 3(2):185–190, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Bratcher:1974:POS

[BB74] T. L. Bratcher and P. Bhalla. On the properties of an optimal selection procedure. Communications in Statistics, 3(2):191–196, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Bratcher:1975:CBP

[BB75] T. L. Bratcher and R. P. Bland. On comparing binomial probabilities from a Bayesian viewpoint. Communications in Statistics, 4(10): 975–985, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Beckman:1973:AMW

[Bec73] R. J. Beckman. An application of multivariate weighing designs. Communications in Statistics, 1(6):561–565, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Behnen:1975:RHT

[Beh75] Konrad Behnen. The Randles–Hogg test and an alternative proposal. Communications in Statistics, 4(3):203–238, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Beaver:1975:MIW

[BG75] R. J. Beaver and D. V. Gokhale. A model to incorporate within-pair order effects in paired comparisons. Communications in Statistics, 4(10):923–939, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Burdick:1974:EMU

[BHOO74] D. S. Burdick, D. G. Herr, W. M. O’Fallon, and B. V. O’Neill. Exact methods in the unbalanced, two-way analysis of variance — a geometric view. Communications in Statistics, 3(6):581–595, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Bagshaw:1975:SDD

[BJ75] Michael Bagshaw and Richard A. Johnson. Sequential detection of a drift change in a Wiener process. Communications in Statistics, 4(8):787–796, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Boardman:1973:EPB

[BK73] Thomas J. Boardman and Peter J. Kendell. Estimators for the parameters of a binomial mixture of grouped-censored exponentials — useful in life testing problems. Communications in Statistics, 2 (1):29–44, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Blischke:1974:NEI

[Bli74] W. R. Blischke. On nonregular estimation. II. Estimation of the location parameter of the gamma and Weibull distributions. Communications in Statistics, 3(12):1109–1129, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Blumenthal:1975:SEL

[Blu75] Saul Blumenthal. Sequential estimation of the largest normal mean when the variance is unknown. Communications in Statistics, 4(7): 655–669, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Basu:1974:EIC

[BOL74] J. P. Basu, P. L. Odell, and T. O. Lewis. The effects of intraclass correlation on certain significance tests when sampling from multivariate normal population. Communications in Statistics, 3(9):899– 908, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Bai:1974:IWT

[BR74a] Do Sun Bai and J. S. Rustagi. Inference in a waiting-time distribution derived from dependent Bernoulli trials. Communications in Statistics, 3(4):361–371, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Brown:1974:JCI

[BR74b] G. H. Brown and D. S. Robson. Joint confidence intervals for linear combinations of, at most, q of p means of normal distributions. Communications in Statistics, 3(3):213–222, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Broemeling:1974:BIA

[Bro74] Lyle David Broemeling. Bayesian inferences about a changing sequence of random variables. Communications in Statistics, 3(3): 243–255, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Bryan:1975:EBE

[BS75] T. O’ Bryan and V. Susarla. An empirical Bayes estimation problem with nonidentical components involving normal distributions. Communications in Statistics, 4(11):1033–1042, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Burr:1973:PGS

[Bur73a] Irving W. Burr. Parameters for a general system of distributions to match a grid of α_{3} and α_{4}. Communications in Statistics, 2(1): 1–21, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Burr:1973:SAR

[Bur73b] Irving W. Burr. Some approximate relations between terms of the hypergeometric, binomial and Poisson distributions. Communications in Statistics, 1(4):297–301, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Burstein:1973:CAB

[Bur73c] Herman Burstein. Close approximations of binomial confidence limits at 10 confidence levels. Communications in Statistics, 2(2):185– 187, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Bellhouse:1974:RC

[BW74] D. R. Bellhouse and M. S. Wolynetz. On the regression curve E(y) = α + δx + βρ^{x}. Communications in Statistics, 3(6):525–534, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Cameron:1975:ERR

[Cam75] M. A. Cameron. An extension of Ruben’s result on the probability integral of noncentral chi-square. Communications in Statistics, 4 (8):783–786, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Carroll:1975:DEU

[Car75] Raymond J. Carroll. Density estimation at unknown points and tail ordering. Communications in Statistics, 4(6):565–574, 1975. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Chhikara:1975:SDR

[CF75] R. S. Chhikara and J. L. Floks. Statistical distributions related to the inverse Gaussian. Communications in Statistics, 4(12):1081– 1091, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Calinski:1974:DMC

[CH74] T. Caliński and J. Harabasz. A dendrite method for cluster analysis. Communications in Statistics, 3(1):1–27, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Chakravorti:1975:MLE

[Cha75] S. R. Chakravorti. Maximum-likelihood estimation for the growth curve model with unique dispersion matrices. Communications in Statistics, 4(1):69–85, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Chen:1974:NRU

[Che74a] C. W. Chen. A note on Rao’s unified theory of least squares. Communications in Statistics, 3(9):909–912, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Chen:1974:OPP

[Che74b] Chao W. Chen. An optimal property of principal components. Communications in Statistics, 3(10):979–983, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Choi:1975:EIO

[Cho75] Jae-Rong Choi. An equality involving orthant probabilities. Communications in Statistics, 4(12):1167–1175, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Cleveland:1974:MDA

[CL74] Williams S. Cleveland and Peter A. Lachenbruch. A measure of divergence among several populations. Communications in Statistics, 3(3):201–211, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Clave:1974:CMA

[Cla74] James T. Mc Clave. A comparison of moving average estimation procedures. Communications in Statistics, 3(9):865–883, 1974. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Consul:1975:CLP

[Con75] P. C. Consul. On a characterization of Lagrangian Poisson and quasi-binomial distributions. Communications in Statistics, 4(6): 555–563, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Cooke:1974:TCS

[Coo74] W. P. Cooke. A test for convexity of a second-order surface. Communications in Statistics, 3(10):1007–1013, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Cleroux:1975:CDS

[CR75] R. Cléroux and P. Robert. The comparison of different scales of measurement in the multivariate normal model: a numerical algorithm. Communications in Statistics, 4(9):839–850, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Crow:1975:CLS

[Cro75] Edwin L. Crow. Confidence limits for small probabilities from Bernoulli trials. Communications in Statistics, 4(5):397–413, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Cassel:1974:ESS

[CS74] Claes-Magnus Cassel and Carl-Erik Sämdal. Evaluation of some sampling strategies for finite populations using a continuous variable framework. Communications in Statistics, 3(4):373–390, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Davidson:1973:RML

[Dav73] Roger R. Davidson. Ranking by maximum likelihood under a model for paired comparisons. Communications in Statistics, 1(5):381– 391, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

David:1974:MLE

[Dav74] David W. David, Jr. Maximum likelihood estimates of the parameters of a mixture of two regression lines. Communications in Statistics, 3(10):995–1006, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Davidson:1974:MTE

[DD74] Roger R. Davidson and Daniel L. Daniel. Moment-type estimation in the exponential family. Communications in Statistics, 3(11):1101– 1108, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Devore:1975:LNN

[Dev75] Jay L. Devore. Linear near-neighbor classifiers for correlated categories. Communications in Statistics, 4(10):891–905, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Dixons:1974:DRL

[DK74] Wilfrid J. Dixons, Ph. D. and Jan W. Kuzma, Ph. D. Data reduction in large clinical trials. Communications in Statistics, 3(4):301–324, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Downing:1975:DDM

[DS75] Daryl J. Downing and John G. Saw. Does the dispersion matrix of a multivariate normal population have Markov structures? Communications in Statistics, 4(11):1073–1079, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Draper:1973:EFA

[DT73] Norman R. Draper and David E. Tierney. Exact formulas for additional terms in some important series expansions. Communications in Statistics, 1(6):495–524, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Dudewicz:1974:EOP

[Dud74] Edward J. Dudewicz. Estimation of ordered parameters: the range of k normal means. Communications in Statistics, 3(3):279–285, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Dunn:1974:MTC

[Dun74] Olive Jean Dunn. On multiple tests and confidence intervals. Communications in Statistics, 3(1):101–103, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

deWaal:1974:BEN

[dW74] D. J. de Waal. Bayes estimate of the noncentrality parameter in multivariate analysis. Communications in Statistics, 3(1):73–79, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Easterling:1975:RSI

[Eas75] Robert G. Easterling. Randomization and statistical inference. Communications in Statistics, 4(8):723–735, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Eastman:1973:PML

[EB73] James Eastman and Lee J. Bain. A property of maximum likelihood estimators in the presence of location-scale nuisance parameters. Communications in Statistics, 2(1):23–28, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Eichhorn:1974:SSO

[Eic74] B. H. Eichhorn. Sequential search of an optimal dosage for cases of linear dosage — toxicity regression. Communications in Statistics, 3(3):263–271, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Farmer:1975:VSM

[FF75] J. H. Farmer and R. J. Freund. Variable selection in the multivariate analysis of variance (MANOVA). Communications in Statistics, 4 (1):87–93, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Funk:1975:CRV

[FR75] G. M. Funk and R. H. Rodine. Characterizing random variables by specifying the distribution of U-shaped functions of their weighted sums. Communications in Statistics, 4(2):143–162, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Friel:1975:APS

[Fri75] Jomes O. Friel. Adaptive prediction of stationary time series by modified conjugate direction methods. Communications in Statistics, 4(1):19–32, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Fu:1974:NUA

[Fu74] James C. Fu. On non-uniqueness of ancillary statistics. Communications in Statistics, 3(11):1093–1100, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gregory:1975:DPU

[GA75] W. C. Gregory and R. L. Anderson. Design procedures and use of prior information in the estimation of parameters of a non-linear model. Communications in Statistics, 4(5):483–496, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Galambos:1975:CTP

[Gal75] Janos Galambos. Characterizations in terms of properties of the smaller of two observations. Communications in Statistics, 4(3): 239–244, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gupta:1975:ADD

[GCK75] K. Gupta, A. K. Chattopadhyay, and R. Krishnaiah. Asymptotic distributions of the determinants of some random matrices. Communications in Statistics, 4(1):33–47, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gray:1974:SAE

[GD74] Henry L. Gray and Benjamin S. Duran. A space application of an extension of the Buffon needle problem. Communications in Statistics, 3(8):807–8, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gelfand:1974:RSM

[Gel74] Alan E. Gelfand. Rapid seriation methods for multivariate observations through similarities. Communications in Statistics, 3(7):635– 645, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gupta:1974:DST

[GG74] R. P. Gupta and D. Gekabe. Distributions of sums of truncated discrete variates with applications. Communications in Statistics, 3 (12):1161–1170, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gupta:1975:DQT

[GG75] A. K. Gupta and Z. Govtndarajulu. Distribution of the quotient of two independent Hotelling’s T^{2}-variates. Communications in Statistics, 4(5):449–453, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gupta:1974:MSS

[GH74] Shanti S. Gupta and Wen-Tao Huang. On a maximin strategy for sampling based on selection procedures from several populations. Communications in Statistics, 3(4):325–359, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Godambe:1975:NAD

[GH75] A. V. Godambe and W. L. Harkness. Normal approximation to the distribution of a cell entry in a 2 × 2 × 2 contingency table. Communications in Statistics, 4(8):699–709, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ghosh:1974:GU

[Gho74] Pankaj Ghosh. On generalized unimodality. Communications in Statistics, 3(6):567–580, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Griffiths:1973:HRM

[GM73] D. A. Griffiths and A. J. Miller. Hyperbolic regression — a model based on two-phase piecewise linear regression with a smooth transition between regimes. Communications in Statistics, 2(6):561–569, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ghosh:1975:SPM

[GP75] Malay Ghosh and Edward Pollak. Some properties of multivariate distributions with pdf’s constant on ellipsoids. Communications in Statistics, 4(12):1157–1160, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Graham:1973:CET

[Gra73] Jack E. Graham. Composite estimation in two cycle rotation sample designs. Communications in Statistics, 1(5):419–431, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gupta:1974:CDP

[Gup74] Ramesh C. Gupta. Characterization of distributions by a property of discrete order statistics. Communications in Statistics, 3(3):287– 289, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gupta:1975:CDC

[Gup75a] Ramesh C. Gupta. On characterization of distribution by conditional expectation. Communications in Statistics, 4(1):99–103, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gupta:1975:SCD

[Gup75b] Ramesh C. Gupta. Some characterizations of discrete distributions by properties of their moment distributions. Communications in Statistics, 4(8):761–765, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gunst:1975:RAP

[GW75] R. F. Gunst and J. T. Webster. Regression analysis and problems of multicollinearity. Communications in Statistics, 4(3):277–292, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Gray:1973:JSR

[GWS73] H. L. Gray, T. A. Watkins, and W. R. Schucany. On the jackknife statistic and its relation to UMVU estimators in the normal case. Communications in Statistics, 2(4):285–320, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Haack:1975:SEB

[Haa75] Dennis G. Haack. Stringency in the empirical Bayes problem. Communications in Statistics, 4(2):185–192, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hamdan:1974:PCS

[Ham74a] M. A. Hamdan. On Pearson’s chi-squared for contingency tables and zero bivariate binomial correlation. Communications in Statistics, 3(9):885–892, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hamdan:1974:UOP

[Ham74b] M. A. Hamdan. The use of orthogonal polynomials and orthonormal functions in the calculation of the noncentrality parameter of chisquared. Communications in Statistics, 3(2):157–166, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hawkins:1977:CNS

[Haw77] Douglas M. Hawkins. Comment on “A new statistic for testing suspected outliers”. Communications in Statistics: Theory and Methods, 6(5):435–438, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Tik75, Tik77].

Hazra:1973:SSW

[Haz73] P. K. Hazra. On the structure of singular weighing designs. Communications in Statistics, 2(6):573–579, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hobbs:1974:AEK

[HC74] G. R. Hobbs and W. J. Conover. The asymptotic efficiency of the Kruskal–Wallis test and the median test in contingency tables with ordered categories. Communications in Statistics, 3(12):1131–1138, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hensler:1975:AEN

[Hen75] Gary L. Hensler. Asymptotically efficient nonparametric estimation of regression and scale parameters. Communications in Statistics, 4(9):821–837, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Higgins:1975:GMC

[Hig75] J. J. Higgins. A geometrical method of constructing multivariate densities and some related inferential procedures. Communications in Statistics, 4(10):955–966, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hoerl:1975:RRS

[HKB75] Arthur E. Hoerl, Robert W. Kannard, and Kent F. Baldwin. Ridge regression: some simulations. Communications in Statistics, 4(2): 105–123, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hong:1973:EAE

[HL73] Dun-Mow Hong and Wilford L. L’Ésperance. Effects of autocorrelated errors on various least squares estimators a Monte Carlo study. Communications in Statistics, 2(6):507–523, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hettmansperger:1973:TSC

[HM73] Thomas P. Hettmansperger and Joseph W. Mckean. On testing for significant chance in c × c tables. Communications in Statistics, 2(6):551–560, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hochberg:1975:SIU

[Hoc75a] Yosef Hochberg. Simultaneous inference under Behrens–Fisher conditions: a two sample approach. Communications in Statistics, 4 (12):1109–1119, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hochberg:1975:TSR

[Hoc75b] Yosef Hochberg. Two stage randomized response schemes for estimating a multinomial. Communications in Statistics, 4(11):1021– 1032, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hutcheson:1974:SME

[HS74] K. Hutcheson and L. R. Shenton. Some moments of an estimate of Shannon’s measure of information. Communications in Statistics, 3 (1):89–94, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hoffman:1975:DLS

[HS75] Thomas R. Hoffman and John G. Saw. Distribution of the largest of a set of equicorrelated normal variables. Communications in Statistics, 4(1):49–53, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Hoal:1975:CTD

[HW75] David G. Hoal and George H. Weiss. A clinical trial design with a fixed maximum number of failures. Communications in Statistics, 4(5):429–436, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Iman:1974:USA

[Ima74] Ronald L. Iman. Use of a t-statistic as an approximation to the exact distribution of the Wilcoxon signed ranks test statistic. Communications in Statistics, 3(8):795–806, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Jain:1975:GNT

[Jai75] C. C. Jain. A generalized Neyman’s Type A contribution for the busy period of queues. Communications in Statistics, 4(11):1065– 1071, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Janardan:1975:CIP

[Jan75] K. G. Janardan. Certain inference problems for multivariate hypergeometric models. Communications in Statistics, 4(4):375–387, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Johnson:1974:PGR

[JH74] Dallas E. Johnson and Victor Hegemann. Procedures to generate random matrices with noncentral distributions. Communications in Statistics, 3(7):691–699, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Johnson:1975:SGF

[JK75] Norman L. Johnson and Samuel Kott. On some generalized Farlie– Gumbel–Morgenstern distributions. Communications in Statistics, 4(5):415–427, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Johnson:1973:SST

[Joh73] N. L. Johnson. Some simple tests of mixtures with symmetrical components. Communications in Statistics, 1(1):17–25, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

John:1974:MUM

[Joh74] S. John. Median-unbiased most acceptable estimates of Poisson, binomial and negative-binomial distributions. Communications in Statistics, 3(12):1155–1159, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kabir:1974:ELS

[KA74a] A. B. M. Lutful Kabir and M. Ahsanullah. Estimation of the location and scale parameters of a power-function distribution by linear functions of order statistics. Communications in Statistics, 3(5):463– 467, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kshirsagar:1974:SSP

[KA74b] Anant M. Kshirsagar and Ersen Arsevan. Stationary state probabilities of a Markov renewal process and optimum scores associated with the states. Communications in Statistics, 3(10):923–931, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kabe:1975:SRG

[Kab75] D. G. Kabe. Some results for the GMANOVA model. Communications in Statistics, 4(9):813–820, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Klimko:1975:TNV

[KAR75] L. A. Klimko, C. E. Ancle, and A. Rademaker. Tests for normality versus lognormality. Communications in Statistics, 4(11):1009–1019, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kastenbaum:1974:ACD

[Kas74] Marvin A. Kastenbaum. Analysis of categorical data: some wellknown analogues and some new concepts. Communications in Statistics, 3(5):401–417, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Katti:1973:EDC

[Kat73] S. K. Katti. Exact distribution for the chi square test in the one way table. Communications in Statistics, 2(5):435–447, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kleyle:1975:EPM

[KD75] Robert M. Kleyle and Ram C. Dahiya. Estimation of parameters of mixed failure time distribution from censored data. Communications in Statistics, 4(9):873–882, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kim:1974:NCP

[Kim74] P. J. Kim. The normal centroid and powers of the Pearson’s χ^{2} test and the Kolmogorov–Smirnov test. Communications in Statistics, 3(3):297–300, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kotz:1974:CED

[KJ74] Samuel Kotz and Norman L. Johnson. A characterization of exponential distributions by a waiting time property. Communications in Statistics, 3(3):257–258, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kemp:1974:FDD

[KK74] Adrienne W. Kemp and C. D. Kemp. A family of discrete distributions defined via their factorial moments. Communications in Statistics, 3(12):1187–1196, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kleinbaum:1973:TLH

[Kle73] David G. Kleinbaum. Testing linear.hypotheses in generalized multivariate linear models. Communications in Statistics, 1(5):433–457, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kaminsky:1975:UMP

[KLN75] Kennetn S. Kaminsky, Eugene M. Luks, and Paul T. Nelson. An urn model and the prediction of order statistics. Communications in Statistics, 4(3):245–250, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kagan:1973:EDS

[KLR73] A. M. Kagan, Yu. V. Linnik, and C. R. Rao. Extension of Darmois– Skitovic theorem to functions of random variables satisfying an addition theorem. Communications in Statistics, 1(5):471–474, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Korwar:1975:CSD

[Kor75] R. M. Korwar. On characterizing some discrete distributions by linear regression. Communications in Statistics, 4(12):1133–1147, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kabir:1974:BEV

[KR74] A. B. M. Lutful Kabir and M. Rahman. Bounds for expected values of order statistics. Communications in Statistics, 3(6):557–566, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Khatri:1974:ELP

[KS74a] C. G. Khatri and K. R. Shah. Estimation of location parameters from two linear models under normality. Communications in Statistics, 3(7):647–663, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kshirsagar:1974:MCD

[KS74b] Anant M. Kshirsagar and Luisa S. Sia. C-matrix of confounded designs in symmetrical factorial experiments. Communications in Statistics, 3(11):1015–1024, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Karson:1975:DCM

[KS75a] Marvin J. Karson and M. Lynn Spruill. Design criteria and minimum bias estimation. Communications in Statistics, 4(4):339–355, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kimeldorf:1975:OPF

[KS75b] George Kimeldorf and Allan Sampson. One-parameter families of bivariate distributions with fixed marginals. Communications in Statistics, 4(3):293–301, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kimeldorf:1975:URB

[KS75c] George Kimeldorf and Allan Sampson. Uniform representations of bivariate distributions. Communications in Statistics, 4(7):617–627, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Kim:1975:UCH

[KV75] Bock Ki Kim and J. Van Ryzin. Uniform consistency of a histogram density estimator and modal estimation. Communications in Statistics, 4(4):303–315, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Krishnaiah:1973:DLC

[KW73] P. R. Krishnaiah and V. B. Walkar. On the distribution of a linear combination correlated quadratic forms. Communications in Statistics, 1(4):371–380, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lepage:1975:AOR

[Lep75] Yves Lepage. Asymptotically optimum rank tests for contiguous location and scale alternatives. Communications in Statistics, 4(7): 671–687, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lurie:1974:GFT

[LHS74] D. Lurie, H. O. Hartley, and M. R. Stroud. A goodness of fit test for censored data. Communications in Statistics, 3(8):745–753, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lin:1974:AME

[Lin74a] Pi-Eih Lin. Admissible minimax estimators of the multivariate normal mean with squared error loss. Communications in Statistics, 3 (1):95–99, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lingaraj:1974:CEC

[Lin74b] B. P. Lingaraj. Certainty equivalent of a chance constraint if the random variable is uniformly distributed. Communications in Statistics, 3(10):949–951, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Littell:1974:RES

[Lit74] Ramon C. Littell. On the relative efficiency of some Kolmogorov– Smirnov-type tests for symmetry about zero. Communications in Statistics, 3(11):1069–1076, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lurie:1973:EIS

[LM73] Dan Lurie and Robert L. Mason. Empirical investigation of several techniques for computer generation of order statistics. Communications in Statistics, 2(4):363–371, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Locke:1975:DXX

[Loc75] Charles Locke. On the dependence of X_{1} + X_{2} and X_{1}/X_{2}. Communications in Statistics, 4(4):357–362, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lott:1973:OSP

[Lot73] William F. Lott. The optimal set of principal component restrictions on a least-squares regression. Communications in Statistics, 2(5): 449–464, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lachenbruch:1973:RLQ

[LSR73] Peter A. Lachenbruch, Cheryl Sneeringer, and Lawrence T. Revo. Robustness of the linear and quadratic discriminant function to certain types of non-normality. Communications in Statistics, 1(1):39– 56, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lwin:1974:LBP

[Lwi74] Thaung Lwin. A lower bound of the posterior variance and its estimate. Communications in Statistics, 3(6):597–604, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Mickey:1974:OLS

[MB74] M. Ray Mickey and Patricia M. Britt. Obtaining linear scores from preference pairs. Communications in Statistics, 3(6):501–511, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

McCune:1975:CFE

[MG75] E. D. McCune and H. L. Gray. On Cornish–Fisher expansions with unknown cumulants. Communications in Statistics, 4(11):1043– 1055, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Mayer:1973:MCO

[MH73] Lawrenca S. Mayer and Arlo D. Hendrickson. A method for constructing an optimal regression design after an initial set of input values has been selected. Communications in Statistics, 2(5):465– 477, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Melnick:1974:FDS

[MM74] Edward L. Melnick and John Moussourakis. Filter design for the seasonal adjustment of a time series. Communications in Statistics, 3(12):1171–1186, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

McWilliams:1974:NED

[MP74] Gerald V. McWilliams and James L. Poirot. A note on effects of digital approximation in linear models. Communications in Statistics, 3(9):893–897, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Mann:1973:MGF

[MSF73] Nancy R. Mann, Ernest M. Scheuer, and Kanneth W. Fertig. A mean goodness-of-fit test for the two-parameter Weibull or extremevalue distribution with unknown parameters. Communications in Statistics, 2(5):383–400, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Mehta:1974:EME

[MT74] J. S. Mehta and P. A. V. B. Swamy Temple. Efficient method of estimating the level of a stationary first-order autoregressive process. Communications in Statistics, 3(1):81–88, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Muirhead:1974:PLL

[Mui74] Robb J. Muirhead. Powers of the largest latent root test of Σ = I. Communications in Statistics, 3(6):513–524, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Mullet:1974:GIA

[MW74] Gary M. Mullet and Victor J. Walsh. A generalized inverse algorithm for analysis of experimental designs in a regression framework. Communications in Statistics, 3(2):179–184, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Melnick:1975:ASM

[MY75] Edward L. Melnick and Uri Yechiali. An application of the simplex method for estimation problems related to contingency tables. Communications in Statistics, 4(12):1121–1132, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Nagarsenker:1975:PPW

[Nag75] B. N. Nagarsenker. Percentage points of Wilk’s L_{vc} criterion. Communications in Statistics, 4(7):629–641, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Naik:1975:SSR

[Nai75] Umesh D. Naik. Some selection rules for comparing p processes with a standard. Communications in Statistics, 4(6):519–535, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Nagase:1973:NSD

[NB73] G. Nagase and K. S. Banerjee. A note on second degree polynomial statistics. Communications in Statistics, 1(6):525–531, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Nagarsenker:1975:EDS

[ND75] B. N. Nagarsenker and M. M. Das. Exact distribution of sphericity criterion in the complex case and its percentage points. Communications in Statistics, 4(4):363–374, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ott:1974:CMW

[OC74] L. Ott and J. A. Cornell. A comparison of methods which utilize the integrated mean square error criterion for constructing response surface designs. Communications in Statistics, 3(11):1053–1068, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ott:1973:WFE

[OM73] L. Ott and W. Mendenhall. Weighted factorial experiments for analysing untransformed data. Communications in Statistics, 2(6): 535–549, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Owen:1973:Ea

[Owe73a] D. B. Owen. Editorial. Communications in Statistics, 1(5):i, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Owen:1973:Eb

[Owe73b] D. B. Owen. Editorial. Communications in Statistics, 1(6):i, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Pagano:1973:WAS

[Pag73] Marcello Pagano. When is an altoregressive scheme stationary. Communications in Statistics, 1(6):533–544, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Patel:1975:GFT

[Pat75] Kantilal M. Patel. A generalized Friedman test for randomized block designs when observations are subject to arbitrary right censorship. Communications in Statistics, 4(4):389–395, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Petrasovits:1974:BAE

[PC74] Andrés Petrásovits and Richard G. Cornell. Bayesian analysis for an exponential surveillance model. Communications in Statistics, 3(7):679–689, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Petrasovits:1975:ABE

[PC75] Andrés Petrásovits and Richard G. Cornell. Approximations to the Bayes estimate for a quantal assay with simple exponential tolerance distribution. Communications in Statistics, 4(9):851–862, 1975. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Perng:1974:FWC

[Per74] S. K. Perng. Fixed width confidence bands for linear regression models. Communications in Statistics, 3(5):495–500, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Paik:1973:PBD

[PF73] U. B. Paik and W. T. Federer. Partially balanced designs act properties A and B. Communications in Statistics, 1(4):331–350, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Patal:1973:GJS

[PH73] Kantilal M. Patal and David G. Hoal. A generalized Jonckheere k-sample test against ordered alternatives when observations are subject to arbitrary right censorship. Communications in Statistics, 2(4):373–380, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Philippeu:1973:EAL

[Phi73] Andreas N. Philippeu. Exponential approximation of the likelihood function and hypotheses testing in the independent not identically distributed case. Communications in Statistics, 2(5):401–413, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Pierce:1975:TA

[Pie75] David A. Pierce. On trend and autocorrelation. Communications in Statistics, 4(2):163–175, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Patal:1973:CSS

[PP73] Jagdish K. Patal and Jagdish K. Patal. Complete sufficient statistics and estimators. Communications in Statistics, 2(4):327–336, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Pillai:1974:MRL

[PP74] K. C. S. Pillai and Dennis L. Young Purdue. On the max U-ratio and likelihood ratio tests of equality of several covariance matrices. Communications in Statistics, 3(1):29–53, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Philippou:1974:NML

[PR74] Andreas N. Philippou and George G. Roussas. A note on the multivariate logarithmic series distribution. Communications in Statistics, 3(5):469–472, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Perlman:1975:SRE

[PR75a] Michael D. Perlman and Uttara A. Rasmussen. Some remarks on estimating a noncentrality parameter. Communications in Statistics, 4(5):455–468, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Pesotan:1975:REA

[PR75b] H. Pesotan and B. L. Raktoe. Remarks on extension of the Anderson–Federer (0, 1)-matrix procedure for main effects to effects of a higher degree. Communications in Statistics, 4(9):797– 811, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Peterson:1974:RMT

[PS74a] R. E. Peterson and K. K. Seo. A regression model test of the overall correctness of financial statements. Communications in Statistics, 3(6):549–556, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Pillai:1974:DHT

[PS74b] K. C. S. Pillai and Sudjana. On the distribution of Hotelling’s trace and power comparisons. Communications in Statistics, 3(5):433– 454, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Pillai:1973:MTR

[PY73] K. C. S. Pillai and Dennis L. Young. The max trace-ratio test of the hypothesis H_{0}: Σ_{−1} = … = Σ_{−k} = λΣ_{−0}. Communications in Statistics, 1(1):57–80, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Quesenberry:1975:TST

[Que75] C. P. Quesenberry. Transforming samples from truncation parameter distributions to uniformity. Communications in Statistics, 4 (12):1149–1155, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Raghunandanan:1974:PWS

[Rag74] K. Raghunandanan. On play the winner sampling rule. Communications in Statistics, 3(8):769–776, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Raktoe:1974:GFU

[Rak74] B. L. Raktoe. A geometrical formulation of an unsolved problem in fractional factorial designs. Communications in Statistics, 3(10): 959–968, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Rao:1973:UTL

[Rao73] C. Radhakrishna Rao. Unified theory of least squares. Communications in Statistics, 1(1):1–8, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Rao:1974:CMN

[Rao74] B. L. S. Prakasa Rao. On a characterization of multivariate normal distribution. Communications in Statistics, 3(11):1077–1080, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Read:1974:CC

[Rea74] Campbell B. Read. Confidence and credibility. Communications in Statistics, 3(3):273–277, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Randles:1973:ADF

[RH73] Ronald H. Randles and Robert V. Hogg. Adaptive distribution-free tests. Communications in Statistics, 2(4):337–356, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Rothman:1975:EVT

[RL75] E. D. Rothman and A. M. Liebetrau. Estimation of the variancetime and covariance-time curves of a stationary point process. Communications in Statistics, 4(4):317–338, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Raghavarao:1974:NMC

[RN74] D. Raghavarao and G. Nageswararao. A note on a method of construction of designs for two-way elimination of heterogeneity. Communications in Statistics, 3(2):197–199, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Robson:1974:PEA

[Rob74] D. S. Robson. On the possibility of exploiting autocorrelations within rows to estimate genetic and environmental variance and covariance in a plantation. Communications in Statistics, 3(10):913– 922, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Rohatgi:1973:LDP

[Roh73] V. K. Rohatgi. On large deviation probabilities for sums of random variables which are attracted to the normal law. Communications in Statistics, 2(6):525–533, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Rao:1974:CSS

[RR74] K. C. Rao and B. S. Robson. A chi-square statistic for goodness-of-fit tests within the exponential family. Communications in Statistics, 3 (12):1139–1153, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ruben:1974:NRP

[Rub74a] Harold Ruben. A new result on the probability integral of noncentral chi-square with even degrees of freedom. Communications in Statistics, 3(5):473–476, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ruben:1974:NCC

[Rub74b] Harold Ruben. Non-central chi-square and gamma revisited. Communications in Statistics, 3(7):607–633, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Rogers:1974:TEW

[RY74] G. S. Rogers and D. L. Young. Testing and estimation when a normal covariance matrix has intraclass structure of arbitrary order. Communications in Statistics, 3(4):343–359, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Rogers:1975:SLR

[RY75] G. S. Rogers and D. L. Young. Some likelihood ratio tests when a normal covariance matrix has certain reducible linear structures. Communications in Statistics, 4(6):537–554, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Ryan:1974:ESC

[Rya74] Thomas A. Ryan, Jr. Estimation from samples containing bad repetitions. Communications in Statistics, 3(4):391–399, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Srivastava:1974:CDM

[SA74] J. N. Srivastava and D. A. Anderson. A comparison of the determinant, maximum root, and trace optimality criteria. Communications in Statistics, 3(10):933–940, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sahai:1974:SFB

[Sah74] Hardeo Sahai. Some formal Bayes estimators of variance components in the balanced three-stage nested random effects model. Communications in Statistics, 3(3):233–242, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Saw:1973:JST

[Saw73] John G. Saw. Jacobians of singular transformations with applications to statistical distribution theory. Communications in Statistics, 1(1):81–91, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Saw:1974:LBD

[Saw74] John G. Saw. A lower bound for the distribution of a partial product of latent roots. Communications in Statistics, 3(7):665–669, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Saw:1975:TIP

[Saw75a] John G. Saw. Tests on the intensity of a Poisson process. Communications in Statistics, 4(8):777–782, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Saw:1975:IAF

[Saw75b] Jonn O. Saw. An improved analytic form for the Jacobian of a singular transformation. Communications in Statistics, 4(3):273– 276, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Stevenson:1974:BSP

[SB74] D. J. Stevenson and R. P. Bland. A Bayes solution to the problem of comparing-tail probabilities of normal populations. Communications in Statistics, 3(12):1197–1204, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Smith:1975:EPL

[SB75a] Robert M. Smith and Lee J. Bain. An exponential power life-testing distribution. Communications in Statistics, 4(5):469–481, 1975. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Susarla:1975:EBT

[SB75b] V. Susarla and T. O’ Bryan. An empirical Bayes two action problem with nonidentical components for a translated exponential distribution. Communications in Statistics, 4(8):767–775, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Samuel-Cahn:1974:RSTa

[SC74a] Ester Samuel-Cahn. Repeated significance test II, for hypotheses about the normal distribution. Communications in Statistics, 3(8): 711–733, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Samuel-Cahn:1974:RSTb

[SC74b] Ester Samuel-Cahn. Repeated significance tests I and II. Generalizations. Communications in Statistics, 3(8):735–744, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Samuel-Cahn:1974:TKR

[SC74c] Ester Samuel-Cahn. Two kinds of repeated significance tests, and their application for the uniform distribution. Communications in Statistics, 3(5):419–431, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Scheaffer:1973:APS

[Sch73a] R. L. Scheaffer. Asymptotic properties of sample proportions under a Markovian renewal structure. Communications in Statistics, 2(2): 167–184, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Scheaffer:1973:TUC

[Sch73b] R. L. Scheaffer. Tests for uniform clustering and randomness. Communications in Statistics, 2(6):479–492, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Simmons:1975:PD

[SD75] D. J. Simmons and J. A. Davis. Pair designs. Communications in Statistics, 4(3):255–272, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sen:1974:EBD

[Sen74] Pranab Kumar Sen. On the estimation of basic distributions in randomized response models. Communications in Statistics, 3(11): 1081–1092, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sinha:1975:DRM

[SG75] Bimal Kumar Sinha and N. Giri. On the distribution of a random matrix. Communications in Statistics, 4(11):1057–1063, 1975. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Shaked:1975:NEG

[Sha75a] Moshe Shaked. A note on the exchangeable generalized Farlie– Gumbel–Morgenstern distributions. Communications in Statistics, 4(8):711–721, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Shapiro:1975:ELS

[Sha75b] S. H. Shapiro. Estimation of location and scale parameters — a compromise. Communications in Statistics, 4(12):1093–1108, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Shuster:1974:VPS

[Shu74] J. J. Shuster. The validation of a petition by a sample survey. Communications in Statistics, 3(3):291–296, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sielkan:1973:SSM

[Sie73] R. L. Sielkan, Jr. Sequential sampling and maximum likelihood estimation. Communications in Statistics, 1(6):567–568, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Singh:1973:PCR

[Sin73] Jagbir Singh. Paired comparison ranking by linear programming. Communications in Statistics, 1(4):351–364, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Singpurwalla:1974:EJP

[Sin74] Nozer D. Singpurwalla. Estimation of the join point in a heteroscedastic regression model arising in accelerated life tests. Communications in Statistics, 3(9):853–863, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Singh:1975:NSD

[Sin75] Jagbir Singh. A note on the Stirling distribution of the second kind. Communications in Statistics, 4(8):753–759, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Shah:1973:EMF

[SK73a] B. K. Shah and C. G. Khatri. An extension to a method of fitting regression curve. Communications in Statistics, 1(4):365–370, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Shah:1973:CRC

[SK73b] K. R. Shah and C. G. Khatri. Connectedness in row-column designs. Communications in Statistics, 2(6):571–573, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Lae:1975:NEM

[sL75] Jack Chao sheng Lae. A note an equal-mean distribution. Communications in Statistics, 4(3):251–254, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sliah:1975:UPI

[Sli75] S. M. Sliah. Use of a priori information in the estimation of a parameter in life testing. Communications in Statistics, 4(2):177–184, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sugiura:1973:AFD

[SN73] Nariaki Sugiura and Hisao Nagao. Asymptotic formulas for the distributions of two trace statistics for testing the independence under fixed alternative. Communications in Statistics, 1(4):285–295, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sugiura:1973:ADM

[SO73] Nariaki Sugiura and Masanori Otake. Approximate distribution of the maximum of C − 1 χ^{2}-statistics (2 × 2) derived from 2 × C contingency table. Communications in Statistics, 1(1):9–16, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sansing:1974:DSN

[SO74a] R. C. Sansing and D. B. Owen. The density of the t-statistic for non-normal distributions. Communications in Statistics, 3(2):139– 155, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sugiura:1974:EMH

[SO74b] Nariaki Sugiura and Masanori Otake. An extension of the Mantel– Haenszel procedure to k 2×C contingency tables and the relation to the logit model. Communications in Statistics, 3(9):829–842, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Scheuren:1975:DAA

[SO75] Frederick J. Scheuren and H. Lock Oh. A data analysis approach to fitting square tables. Communications in Statistics, 4(7):595– 615, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Searle:1973:WIM

[SR73] S. R. Searle and J. W. Rudan. Wanted: an inverse matrix. Communications in Statistics, 2(2):155–166, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sedransk:1974:SSA

[SS74] J. Sedransk and Bahadur Singh. Sample size allocation in twophase sampling. Communications in Statistics, 3(11):1025–1040, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Subrahmaniam:1975:CRC

[SS75] Kathleen Subrahmaniam and Kocherlakota Subrahmaniam. On the confidence region comparison of some solutions for the multivariate Behrens–Fisher problem. Communications in Statistics, 4(1):57– 67, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Srivastava:1975:SOM

[ST75] V. K. Srivastava and Ramji Tiwari. The second order moment matrix of generalized double k-class estimators. Communications in Statistics, 4(6):575–589, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Steck:1974:ESS

[Ste74] G. P. Steck. Evaluation of some Steck determinants with applications. Communications in Statistics, 3(2):121–138, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sugiura:1973:DCR

[Sug73] Nariaki Sugiura. Derivatives of the characteristic root of a symmetric or a Hermitian matrix with two applications in multivariate analysis. Communications in Statistics, 1(5):393–417, 1973. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sugiura:1974:MLE

[Sug74] Nariaki Sugiura. Maximum likelihood estimates for the logit model and the iterative scaling method. Communications in Statistics, 3 (10):985–993, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Sundberg:1974:ETF

[Sun74] Rolf Sundberg. On estimation and testing for the folded normal distribution. Communications in Statistics, 3(1):55–72, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Tan:1973:FPN

[Tan73] Peter Tan. On Fisher’s problem of the Nile in uncertain inference. Communications in Statistics, 2(1):45–58, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Tan:1974:NML

[Tan74] Peter Tan. A note on maximum likelihood estimators and nuisance parameters. Communications in Statistics, 3(10):969–977, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Thomas:1974:TLD

[TC74] Marlin A. Thomas and John R. Crigler. Tolerance limits for the pdimensional radial error distribution. Communications in Statistics, 3(5):477–483, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Thompson:1973:FPF

[Tho73] W. A. Thompson, Jr. Finite populations; foundation for inference? Communications in Statistics, 1(1):27–38, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Tiku:1974:NSTa

[Tik74a] M. L. Tiku. A new statistic for testing for normality. Communications in Statistics, 3(3):223–232, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Tiku:1974:TNE

[Tik74b] M. L. Tiku. Testing normality and exponentiality in multi-sample situations. Communications in Statistics, 3(8):777–794, 1974. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Tiku:1975:NST

[Tik75] M. L. Tiku. A new statistic for testing suspected outliers. Communications in Statistics, 4(8):737–752, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic). See comment [Haw77].

Tiku:1977:RCN

[Tik77] M. L. Tiku. Rejoinder: “Comment on ‘A new Statistic for Testing Suspected Outliers’ ”. Communications in Statistics: Theory and Methods, 6(14):1417–1422, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Tik75, Haw77].

Tubbs:1975:NAM

[TLD75] J. D. Tubbs, T. O. Lewis, and B. S. Duran. A note on the analysis of the MANOVA model and its application to growth curves. Communications in Statistics, 4(7):643–653, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Thomas:1973:ODI

[TM73] Marlin A. Thomas and Raymond H. Myers. Optical designs for the inverse regression method of calibration. Communications in Statistics, 2(5):419–433, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Tithayasai:1974:BS

[TR74] C. Tithayasai and D. S. Robson. A blocking strategy. Communications in Statistics, 3(11):1041–1052, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Tiku:1974:NSTb

[TRM74] M. L. Tiku, K. Rai, and E. Mead. A new statistic for testing exponentiality. Communications in Statistics, 3(5):485–493, 1974. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Tan:1974:EAF

[TS74] Peter Tan and Aly Sherif. An extension and application of Fisher’s bivariate exponential model. Communications in Statistics, 3(5): 455–462, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Urquhart:1973:EAL

[UWH73] N. S. Urquhart, D. L. Weeks, and C. R. Henderson. Estimation associated with linear models: a revisitation. Communications in Statistics, 1(4):303–330, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

VanRyzin:1973:HMD

[Van73] J. Van Ryzin. A histogram method of density estimation. Communications in Statistics, 2(6):493–506, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Vinod:1973:GDW

[Vin73] H. D. Vinod. “Generalization of the Durbin–Watson statistic for higher order autoregressive processes”. Communications in Statistics, 2(2):115–144, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Vithayasai:1973:OTT

[VR73] C. Vithayasai and D. S. Robson. One-and two-tailed simultaneous orthogonal contrasts. Communications in Statistics, 1(5):459–470, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Wang:1973:SES

[Wan73] Y. H. Wang. Sequential estimation of the scale parameter of the Pareto distribution. Communications in Statistics, 2(2):145–154, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Warde:1974:SLR

[War74] William D. Warde. Simple linear regression with inequality constraints on the parameters: a Bayesian approach. Communications in Statistics, 3(6):535–548, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Woolson:1974:CMC

[WC74] Robert F. Woolson and James W. L. Cole. Comparing means of correlated variates with missing data. Communications in Statistics, 3(10):94, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Weiss:1973:SPB

[Wei73] Lionel Weiss. Statistical procedures based on a gradually increasing number of order statistics. Communications in Statistics, 2(2):95– 114, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Woodward:1975:MVU

[WG75] W. A. Woodward and H. L. Gray. Minimum variance unbiased estimation in the gamma distribution. Communications in Statistics, 4(10):907–922, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Watkins:1973:UER

[WK73] T. A. Watkins and D. M. Karn. On UMVU estimators related to the multivariate normal distribution. Communications in Statistics, 2(4):321–325, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Wolfe:1973:THA

[Wol73] Douglas A. Wolfe. On testing hypotheses about function-free birameters. Communications in Statistics, 1(6):475–493, 1973. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Wood:1974:DTC

[Woo74] J. T. Wood. The distribution of transformed concordance coefficients. Communications in Statistics, 3(7):701–710, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Woolson:1974:ACC

[WS74] Robert F. Woolson and Pranab Kumar Sen. Asymptotic comparison of a class of multivariate multiparameter tests. Communications in Statistics, 3(9):813–828, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Wahba:1975:CAF

[WW75a] G. Wahba and S. Wold. A completely automatic French curve: fitting spline functions by cross validation. Communications in Statistics, 4(1):1–17, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Wahba:1975:PSS

[WW75b] G. Wahba and S. Wold. Periodic splines for spectral density estimation: the use of cross validation for determining the degree of smoothing. Communications in Statistics, 4(2):125–141, 1975. CO-DEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Yancey:1974:MSE

[YJB74] T. A. Yancey, G. G. Judge, and M. E. Bock. A mean square error test when stochastic restrictions are used in regression. Communications in Statistics, 3(8):755–768, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).

Zacks:1974:PCE

[Zac74] S. Zacks. The proportional closeness and the expected sample size of sequential procedures for estimating tail probabilities in exponential distributions. Communications in Statistics, 3(2):105–120, 1974. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic).